The table below shows the frequency of outcomes when two distinguishable coins are tossed 4,000 times and the uppermost faces were observed. What is the estimated probability that the second coin lands with tails up? Outcome HH HT TH TT Frequency 1200 1200 600 1000 Round your answer to four decimal places, if necessary.
Answer #1

Table:

Outcome HH HT TH TT Total   
Frequency 1200 1200 600 1000 4000
Probability 1200/4000 = 0.30 =1200/4000 = 0.30 =600/4000 = 0.15 =1000/4000= 0.25 Total = 1

Numer of outcomes of second coin lands with tails = 1200 + 1000 = 2200

Therefore,

Probability that the second coin land with tails up = 2200 / 4000 = 0.5500

OR 0.30 + 0.25 = 0.5500